OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3?

OpenStudy (allank):

write the equation in point-slope form i.e. like y=m(x-x1)+y1 so in point-slope form, equation is y=-3(x-4)+(-2) y=-3x+12-2 y=-3x+10 Then rewrite in standard form: 3x+7=10

OpenStudy (allank):

note that in point-slope form, m=gradient, x1 and y1 are coordinates of a point on the line, in this case 4 and -2

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

dont you have to do something else to seven?? :?

OpenStudy (allank):

Oh yeah, sorry, we had y=-3x+10 So in standard form it should be 3x+y=10 sorry :)

OpenStudy (anonymous):

oh O.O! OK :) THANKS YOU !

OpenStudy (allank):

you're welcome. :)